# NPHardEval: Dynamic Benchmark on Reasoning Ability of Large Language Models via Complexity Classes

### 12 Jan 2024 v3

@@ 1 Introduction
Another novel feature of our benchmark is its end-to-end automation, encompassing both the generation of tasks and the verification of results. This automation is facilitated by the use of well-known tasks within the benchmark, for which mature and established algorithms have been developed to provide solutions. This systematic approach ensures a high degree of accuracy and reliability in the evaluation process, which also enables easy update of datapoints in the benchmark. This automated framework facilitates an effortless updating of datapoints within the benchmark. As a result, we design the benchmark to refresh its datapoints monthly, effectively reducing the likelihood of the model overfitting the dataset. This dynamic updating mechanism is crucial in maintaining the rigor and relevance of the benchmark over time.
@@
In our research utilizing the benchmark, we aim to address three critical questions to evaluate and understand the reasoning abilities of LLMs:
@@
    Reasoning Ability Gap among Foundation Models: How far away are current open-sourced models compared with closed-sourced models? We aim to quantitatively analyze the differences in performance between closed-source models (including as GPT 4 Turbo, Claude 2, GPT 3.5 Turbo, Claude Instant, and PaLM 2) and open-source models (including Vicuna-13b, Yi-34b, Mistral-7b, MPT-30b, and Phi-1.5). This comparison will shed light on the relative strengths and weaknesses of these models in complex reasoning scenarios.
@@
    Reasoning Ability of Foundation Models: How well can LLMs perform on complex reasoning tasks in different complexity classes with different difficulty levels? These tasks are systematically classified into three complexity classes: P, NP-complete, and NP-hard. Further, we delve into the nuances of these categories by introducing gradation, dividing each task into 10 distinct levels of difficulty. The objective here is to determine the proficiency of LLMs in solving progressively challenging problems, thereby gauging their capability to handle tasks with escalating complexity.
@@ 3.1.3 NP-hard problems
These problems are at least as hard as the hardest problems in NP. They may not necessarily be in NP (i.e., they may not have solutions verifiable in polynomial time) but solving an NP-hard problem in polynomial time would imply that P = NP. We include three NP-hard problems that are not reducible to NP-complete problems in the benchmark, namely Traveling Salesman Problem Optimization Version (TSP-O), Graph Coloring Problem Optimization Version (GCP-O), Longest Path Problem Optimization Version (LPP-O), and Meeting Scheduling Problem (MSP).
Traveling Salesman Problem (Optimization Version, TSP-O)
@@
Graph Coloring Problem (Optimization Version, GCP-O)
@@ 4 Experimental Setting
4.1 Experiment 1: Comparative Analysis of Model Performance
@@
4.2 Experiment 2: Model Performance across Task Complexity
In this experiment, our objective was to assess task complexity and problem difficulty using the Wilcoxon test. This non-parametric statistical hypothesis test allows us to compare two populations with matched samples.
To evaluate task complexity, specifically comparing the complexity among P, NP-Complete, and NP-Hard pairs, we initially pinned the data based on complexity levels. Subsequently, we applied the Wilcoxon test to each pair of complexity sets. To evaluate problem difficulty, aiming to discern differences among problems within the complexity category, we pinned the data based on the specific problems and then used the Wilcoxon test to compare pairs of different problem sets.
4.3 Experiment 3: Sanity Check of Finetuning’s Effect on Benchmark Robustness
The primary objective of this experiment is to ascertain whether it is possible to “hack” our benchmark by finetuning models on its previous versions. To simulate this, we constructed five versions of the benchmark, maintaining a consistent difficulty level. Additionally, we utilize two distinct versions of the benchmark, each varying in difficulty, to evaluate the potential for hacking under varying conditions. To replicate the progression of time, models were finetuned sequentially on one to five benchmarks, each finetuned checkpoint is tested on the three distinct benchmarks for evaluation.
@@ 4.3 Experiment 3: Sanity Check of Finetuning’s Effect on Benchmark Robustness
4.4 Experiment 4: Comparative Analysis of Learnability by In-context Learning
@@
  • If a model exhibits the ability to generalize only from more complex problems but is unable to extend this learning to simpler ones, it reveals a deficiency in its capacity for generalization in terms of reasoning. This suggests that the model is not genuinely acquiring problem-solving skills from the examples but merely recognizing and applying patterns from examples that are of equal or greater complexity to the problem at hand.
@@ 5.1 Reasoning Ability of Foundation Models
Experiment 1 focuses on a comprehensive comparison among various foundation models. In Figure 2, we present the overall zero-shot accuracy for each problem, providing a visual representation of the performance of different models. Our observations reveal that closed-source models generally demonstrate higher accuracy and a lower rate of failure compared to their open-source counterparts. Notably, GPT-4 Turbo often emerges as the frontrunner in performance across the majority of tasks, indicating its superior problem-solving capabilities, while Claude 2, on the other hand, often performs the best on medium-level (NP-complete) complexity in zero-shot settings. Within the realm of open-source models, Yi-34b, Qwen-14b, and Mistral-7b distinguish themselves by significantly outperforming other models in this category. This disparity highlights a notable performance gap and suggests that these two models possess more advanced reasoning abilities compared to other open-source options.
@@
In particular, we use the weighted accuracy and the failure rate metrics to further quantify different models’ performance. The trends observed below in both weighted accuracy and failure rates point to a nuanced understanding of the capabilities and limitations of current LLMs.
@@
5.2 Reasoning Performance across Task Complexity and Difficulty Levels
Experiment 2 is designed to assess how the performance of various LLMs varies across different levels of task complexity and question difficulty. Figure 4 shows the accuracy of each model across different complexity levels. The test results reveal statistical significance ($p<0.05$) in the p-values between P and NP-Hard, as well as NP-Complete and NP-Hard. These findings indicate that our investigated LLMs performed significantly worse when confronted with NP-Hard problems compared to P and NP-Complete problems.
Figure 4: Model’s performance on each complexity level. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Qwen-14b. (h) Mistral-7b. (i) Phi-2. (j) MPT-30b. (k) Vicuna-13b. (l) Phi-1.5.
Figure 5: Model’s performance on tasks across complexity levels. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Qwen-14b. (h) Mistral-7b. (i) Phi-2. (j) MPT-30b. (k) Vicuna-13b. (l) Phi-1.5.
@@ 5.2 Reasoning Performance across Task Complexity and Difficulty Levels
Considering NP-Hard complexity as the most intricate task set among the three (as evidenced in Figure 4), many of the examined models encountered challenges in identifying tasks within this complexity. For the GCP task, Mistral-7b, PaLM 2, GPT 3.5 Turbo, and GPT 4 Turbo exhibited some potential, while Vicuna-13b and Claude Instant 1.2 showed limited performance. For the TSP task, identification was observed only in Claude 2 and GPT 4 Turbo. Of all the investigated models, GPT 4 Turbo exhibited promise in identifying these three tasks within the NP-Hard complexity. However, the performance in GCP and TSP identification significantly surpassed that of the MSP task across these models. For the MSP task, only GPT 4 Turbo displayed some ability for identification, albeit with notably low accuracy.
5.3 Evaluating Benchmark Vulnerability to Finetuning Strategies
@@
In this study, we explore the robustness of benchmark against hacking attempts through a process of finetuning on pairs of question and gold answer. We experiment using 3 well-performing open-source models: Qwen-14b, Mistral-7b, and Phi-2 on two versions of benchmarks. Specifically, we present their performance on two algorithmic questions: the polynomial-time solvable SPP problem and the NP-complete KSP Problem.
Figure 6 presents four graphs illustrating these scenarios: (1) the upper left graph shows SPP challenges at difficulty levels 1-10; (2) the lower left graph displays SPP at levels 11-20; (3) the upper right graph represents KSP at levels 1-10; and (4) the lower right graph shows KSP at levels 11-20. The first row of indicate the accuracy mean of each model, which is averaged over outcomes at 5 finetuning checkpoints, ranging from zero (no finetuning) to five, all within difficulty levels 1-10.
Our findings are twofold: (1) While finetuning yields improvements in solving SPP problems, its impact on the more complex NP-complete KSP is minimal. This suggests the inherent difficulty of hacking NP-complete, and potentially NP-hard, problems through basic question-and-answer approaches. Manual annotation of the thought process, which is not provided in the benchmarks, could potentially enhance effectiveness, albeit with challenges in annotation. (2) Finetuning appears beneficial for performance within the same difficulty level for P problems, yet shows limited out-of-distribution (OOD) adaptability and struggles to generalize to more challenging problems (as evidenced in graphs a and c). For instance, Qwen-14b demonstrates notable proficiency on SPP challenges at levels 1-10 following finetuning; its performance is comparable to that of GPT-4. However, its efficacy significantly diminishes on SPP problems at levels 11-20, even underperforming compared to its unfinetuned state. This indicates that finetuning on these benchmarks could actually impede generalization capabilities and increase the complexity of hacking.
In conclusion, our benchmarks are challenging to hack due to two primary factors: (1) the inherent complexity of NP-complete and NP-hard problems, which are difficult to learn solely from question-and-answer pairs, and (2) the propensity for P problems to become overfitted through finetuning on these pairs, which can be easily exposed when increasing the problem difficulty. Based on these conclusions, we intend to periodically update our benchmarks strategetically with varying difficulty levels to minimize the potential for hacking.
@@ 5.4 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
In Experiment 4, we focused on the tasks of SAS and EDP to investigate the nature of the in-context learning capabilities of LLMs. This experiment empirically distinguishes between “learning” and “mimicking” as exhibited by LLMs during in-context learning scenarios. Our findings also revealed a clear dichotomy in the approach to learning and generalization from examples between closed-source and open-source models.
@@ 7 Acknowledgement
We extend our sincere gratitude to Libby Hemphill for her invaluable support in this work. Additionally, we are grateful for the diverse online feedback we received, which illuminated our path towards enhancing the quality of this paper.

### 12 Feb 2024 v4

@@ 1 Introduction
Another novel feature of our benchmark is its end-to-end automation, encompassing both the generation of tasks and the verification of results. This automation is facilitated by the use of well-known tasks within the benchmark, for which mature and established algorithms have been developed to provide solutions. This systematic approach ensures a high degree of accuracy and reliability in the evaluation process, which also enables easy update of datapoints in the benchmark. This automated framework facilitates an effortless updating of datapoints within the benchmark. As a result, we design the benchmark to refresh its datapoints monthly, effectively reducing the likelihood of the model overfitting the dataset. This dynamic updating mechanism is crucial in maintaining the rigor and relevance of the benchmark over time. We welcome open submissions of model performance to our benchmark directly through the NPHardEval leaderboard on HuggingFace: https://huggingface.co/spaces/NPHardEval/NPHardEval-leaderboard.
@@
In our research utilizing the benchmark, we aim to address three critical aspects to evaluate and understand the reasoning abilities of LLMs (Foundation Models):
@@
    Model Performance Comparison: Our benchmark compares the reasoning ability of 12 closed-source models (including as GPT 4 Turbo, Claude 2, GPT 3.5 Turbo, Claude Instant, and PaLM 2) and open-source models (including Yi-34b, Qwen-14b, Mistral-7b, Phi-2, MPT-30b, Vicuna-13b, and Phi-1.5) across three complexity classes (P, NP-complete, and NP-hard) and 10 difficulty levels. This comparison can shed light on the relative strengths and weaknesses of these models and determine the proficiency of them in solving progressively challenging problems, thereby gauging their capability to handle tasks with escalating complexity.
@@ 3.1.3 NP-hard problems
These problems are at least as hard as the hardest problems in NP. They may not necessarily be in NP (i.e., they may not have solutions verifiable in polynomial time) but solving an NP-hard problem in polynomial time would imply that P = NP. We include three NP-hard problems that are not reducible to NP-complete problems in the benchmark, namely Traveling Salesman Problem Optimization Version (TSP), Graph Coloring Problem Optimization Version (GCP), and Meeting Scheduling Problem (MSP).
Traveling Salesman Problem (Optimization Version, TSP)
@@
Graph Coloring Problem (Optimization Version, GCP)
@@ 4 Experimental Setting
4.1 Experiment 1: Model Performance Comparison
@@
To evaluate across task complexity, specifically comparing the complexity among P, NP-Complete, and NP-Hard pairs, we initially pinned the data based on complexity levels. Subsequently, we applied the Wilcoxon test to each pair of complexity sets. Wilcoxon is a non-parametric statistical hypothesis test that allows us to compare two populations with matched samples. To evaluate problem difficulty, aiming to discern differences among problems within the complexity category, we pinned the data based on the specific problems and then used the Wilcoxon test to compare pairs of different problem sets.
4.2 Experiment 2: Benchmark Robustness
The primary objective of this experiment is to ascertain whether it is possible to “hack” our benchmark by finetuning models on its previous versions. To simulate this, we constructed five versions of the benchmark, maintaining a consistent difficulty level. Additionally, we utilize two distinct versions of the benchmark, each varying in difficulty, to evaluate the potential for hacking under varying conditions. To replicate the progression of time, models were finetuned sequentially on one to five benchmarks, each finetuned checkpoint is tested on the two distinct benchmarks for evaluation.
@@ 4.2 Experiment 2: Benchmark Robustness
4.3 Experiment 3: Comparative Analysis of Learnability by In-context Learning
@@
  • If a model exhibits the ability to generalize only from some types of examples but is unable to extend this learning to others, it reveals a deficiency in its capacity for generalization in terms of reasoning. This suggests that the model is not genuinely acquiring problem-solving skills from the examples but merely recognizing and applying patterns from examples that are of equal or greater complexity to the problem at hand.
@@ 5.1 Reasoning Ability of Foundation Models
Experiment 1 focuses on a comprehensive comparison among various foundation models and across complexity classes and difficulty levels. In Figure 2, we present the overall zero-shot accuracy for each problem, providing a visual representation of the performance of different models.
@@
Our observations reveal that closed-source models generally demonstrate higher accuracy and a lower rate of failure compared to their open-source counterparts. Notably, GPT-4 Turbo often emerges as the frontrunner in performance across the majority of tasks, indicating its superior problem-solving capabilities, while Claude 2, on the other hand, often performs the best on medium-level (NP-complete) complexity in zero-shot settings.
Within the realm of open-source models, Yi-34b, Qwen-14b, and Mistral-7b distinguish themselves by significantly outperforming other models in this category. We observe a disparity between the performance of these three models and other open-source options, highlighting a notable performance gap and suggesting that these models possess more advanced reasoning abilitie.
In particular, we use the weighted accuracy and the failure rate metrics to further quantify different models’ performance. The trends observed below in both weighted accuracy and failure rates point to a nuanced understanding of the capabilities and limitations of current LLMs. We also utilize performance comparison tests within and across complexity classes, to further explore the model performance differences among complexity classes.
@@
5.1.1 Performance across Task Complexity and Difficulty Levels
Figure 4 shows the accuracy of each model across different complexity levels. The test results reveal statistical significance ($p<0.05$) in the p-values between P and NP-Hard, as well as NP-Complete and NP-Hard. These findings indicate that our investigated LLMs performed significantly worse when confronted with NP-Hard problems compared to P and NP-Complete problems.
Figure 4: Models’ performance on each complexity level. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Qwen-14b. (h) Mistral-7b. (i) Phi-2. (j) MPT-30b. (k) Vicuna-13b. (l) Phi-1.5.
Figure 5: Models’ performance on tasks across complexity levels. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Qwen-14b. (h) Mistral-7b. (i) Phi-2. (j) MPT-30b. (k) Vicuna-13b. (l) Phi-1.5.
@@ 5.1.1 Performance across Task Complexity and Difficulty Levels
Considering NP-Hard complexity as the most intricate task set among the three (as evidenced in Figure 4), many of the examined models encountered challenges in identifying tasks within this complexity. For the GCP task, Mistral-7b, PaLM 2, GPT 3.5 Turbo, and GPT 4 Turbo exhibited some potential, while Vicuna-13b and Claude Instant 1.2 showed limited performance. For the TSP task, identification was observed only in Claude 2 and GPT 4 Turbo. Of all the investigated models, GPT 4 Turbo exhibited promise in identifying these three tasks within the NP-Hard complexity. However, the performance in GCP and TSP identification significantly surpassed that of the MSP task across these models. For the MSP task, only GPT 4 Turbo displayed some ability for identification, while with notably low accuracy.
5.2 Evaluating Benchmark Robustness
@@
In Experiment 2, we explore the robustness of benchmark against hacking attempts through a process of finetuning on pairs of question and gold answer. We experiment using 3 well-performing open-source models: Qwen-14b, Mistral-7b, and Phi-2 on two versions of benchmarks. Figure 6 presents the result [^2] We do not present the result on MSP as this problem does not have a fixed solution and we do not conduct finetuning on it.: each problem has two graph with one displaying evaluation results at difficulty levels 1-10 and one displaying evaluation results at difficulty levels 11-20. In each graph, the first row of indicate the accuracy mean of each model, averaged over outcomes at 5 finetuning checkpoints, ranging from tuning using zero (no finetuning) to five distinct benchmarks.
Our findings are twofold: (1) While finetuning yields improvements in solving polynomial-time problems, its impact on the more complex NP-complete and NP-hard problems are negative. This suggests the inherent difficulty of hacking NP-complete, and potentially NP-hard, problems through the basic finetuning with question-and-answer approach. Manual annotation of the chain-of-thought, which is not provided in the benchmarks, could potentially enhance effectiveness, albeit with challenges in annotation. (2) Finetuning appears beneficial for performance within the same difficulty level all P problems, yet shows limited out-of-distribution (OOD) adaptability and struggles to generalize to more difficult problems (as evidenced in graphs a and c) except SAS. For instance, Qwen-14b demonstrates notable proficiency on SPP challenges at levels 1-10 following finetuning; its performance is comparable to that of GPT-4. However, its performance significantly diminishes on SPP problems at levels 11-20, even underperforming compared to its unfinetuned checkpoint. This indicates that finetuning on these benchmarks can only benefit very simple questions such as SAS but could potentially impede generalization capabilities and renders finetuning hacking useless.
In conclusion, our benchmark is challenging to hack due to two primary factors: (1) the inherent complexity of NP-complete and NP-hard problems, which are difficult to learn solely from question-answer pairs, and (2) the propensity for P problems to become overfitted through finetuning on these pairs, while the real “reasoning” ability can be easily exposed by increasing the problem difficulty level. Based on these conclusions, we intend to periodically update our benchmarks strategetically with varying difficulty levels to minimize the potential for hacking.
@@ 5.3 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
In Experiment 3, we focused on the tasks of SAS and EDP to investigate the nature of the in-context learning capabilities of LLMs. This experiment empirically distinguishes between “learning” and “mimicking” as exhibited by LLMs during in-context learning scenarios. Our findings also revealed a clear dichotomy in the approach to learning and generalization from examples between closed-source and open-source models.
@@ 7 Acknowledgement
We extend our sincere gratitude to Libby Hemphill for her invaluable support in this work. We also thank Jinkui Chi, who kindly contributed to the maintenance of the benchmark’s code repository and user guidance. Additionally, we are grateful for the diverse feedback we received from Siqi Liu and many others, which illuminated our path towards enhancing the quality of this paper.